# Beyond Linear Approximations: A Novel Pruning Approach for Attention Matrix

### v1

@@ C.2 Matrix Form of $\frac{\mathrm{d}}{\mathrm{d}M}{\cal L}(M)$
  • Let ${\cal L}(M)$ be defined in Definition B.7.
  • Let $p$ be defined in Definition C.1.
@@
where the first step follows from Lemma B.15, the second step follows from Lemma C.2, Lemma C.3, and Lemma C.4, the third step follows from basic algebra, and the fourth step follows from Definition C.1. ∎
Appendix D Bounds for Basic Functions
D.1 Basic Assumptions
Here we introduce our bounded parameters assumption.
Assumption D.1 (Bounded parameters).
@@ D.1 Basic Assumptions
Here we present the lemma of bounds for $M$ and $M_{c}$.
Lemma D.2 (Bounds for $M$ and $M_{c}$).
@@
This Lemma simply follows from the definition of Frobenius norm, given that the max value of each entry in $M$ and $M_{c}$ is $1$. ∎
D.2 Bounds for Basic Functions
We first introduce the lemma of bounds for basic function.
Lemma D.3.
Under Assumption D.1, for all $i_{0}\in[n]$, $j_{0}\in[n]$, $i_{1}\in[d]$, $j_{1}\in[d]$, we have the following bounds
@@ D.2 Bounds for Basic Functions
where the first step follows from Definition B.4, the second step follows from triangle inequality.
@@
Proof of Part 5. The proof simply follows from Assumption D.1 and Fact A.5.
Proof of Part 6. The proof simply follows from Assumption D.1 and Fact A.5.
@@
where the first step follows from Fact A.4 the second step follows from Fact A.5, the third step follows from Part 3., and the last step follows from simple algebra. ∎
D.3 Bounds for Gradient of $\widetilde{f}(M)$
We introduce the lemma of bounds for gradient of $\widetilde{f}(M)$.
Lemma D.4.
@@ D.3 Bounds for Gradient of $\widetilde{f}(M)$
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
  • Assumption D.1 holds.
@@
Appendix E Lipschitz of Gradient
E.1 Useful Facts
Here we introduce the fact of mean value theorem for matrix function.
Fact E.1 (Mean value theorem for matrix function, Fact C.6 in [61]).
@@ E.1 Useful Facts
where the second step follows from Eq. (8), the third step follows from chain rule, the fourth step follows from Cauchy-Schwartz inequality.
@@
Here we introduce the fact of Lipschitz for product of functions.
Fact E.2 (Lipschitz for product of functions, Fact H.2 in [22]).
@@
  • Let $\{f_{i}(x)\}_{i=1}^{n}$ be a sequence of function with same domain and range.
@@ E.2 Lipschitz of $\widetilde{f}(M)$
Lemma E.3 (Lipschitz of $\widetilde{f}(M)$).
@@
  • Assumption D.1 holds.
  • Let $\widetilde{f}(M)$ be defined as Definition B.3.
@@
where the first step follows from Fact E.1, the second step follows from Lemma D.4. ∎
E.3 Lipschitz of $c(M)$
@@
  • Assumption D.1 holds.
  • Let $c(M)$ be defined as Definition B.4.
@@
where the first step follows from Fact E.1, the second step follows from Lemma B.12, the third step follows from Lemma D.4. ∎
E.4 Lipschitz of $\widetilde{f}(M)\circ c(M)$
@@
Lemma E.5 (Lipschitz of $\widetilde{f}(M)\circ c(M)$).
@@
  • Assumption D.1 holds.
  • Let $c(M)$ be defined as Definition B.4.
  • Let $\widetilde{f}(M)$ be defined as Definition B.3.
@@
where the first step follows from triangle inequality, the second step follows from Fact A.5, the third step follows from Lemma D.3, the fourth step follows from Lemma E.4 and Lemma E.3.
@@
E.5 Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})$
@@
Lemma E.6 (Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})$).
@@
  • Assumption D.1 holds.
  • Let $c(M)$ be defined as Definition B.4.
  • Let $\widetilde{f}(M)$ be defined as Definition B.3.
@@
where the first step follows from Fact A.4, the second step follows from basic algebra, the third step follows from Fact A.5, and the fourth step follows from $\|{\bf 1}_{n}\|_{2}=\sqrt{n}$.
@@
which follows from Fact A.5.
Following Eq. (E.5), Eq. (10) and Lemma E.5, we have
@@
E.6 Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})\widetilde{f}(M)$
@@
Lemma E.7 (Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})\widetilde{f}(M)$).
@@
  • Assumption D.1 holds.
  • Let $c(M)$ be defined as Definition B.4.
  • Let $\widetilde{f}(M)$ be defined as Definition B.3.
@@
Following Fact E.2, we have
@@
where we have the upper bound in Lemma D.3, the Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))$ and $\widetilde{f}(M)$ in Lemma E.3 and Lemma E.6. ∎
E.7 Lipschitz of Gradient
@@
Theorem E.8 (Lipschitz of the gradient, formal version of Theorem 5.4).
@@
  • Assumption D.1 holds.
  • Let $c(M)$ be defined as Definition B.4.
  • Let $\widetilde{f}(M)$ be defined as Definition B.3.
@@
where the first step follows from Theorem C.5, and the second step follows from triangle inequality. Now we proof these two terms separately.
@@
where the first step and the second step follows from Fact A.5, the third step follows from triangle inequality, and the fourth step follows from Assumption D.1.
@@
where the first step follows from Lemma E.5 and Lemma E.7, the second step follows from $n\geq 1$.
@@
which follows from Fact A.5.
@@
which follows from Eq. (E.7), Eq. (E.7), and Eq. (13). ∎
Appendix F Convergence of Gradient Descent
F.1 Helpful Statements
Here, we present useful fact that we use to prove our convergence result.
Fact F.1.
@@ F.1 Helpful Statements
F.2 Lower Bound on Frobenius Norm
We present the lemma for the lower bound on the Frobenius norm in this section.
Lemma F.2.
@@ F.2 Lower Bound on Frobenius Norm
where the first step follows from Fact A.6, the second step follows from Fact A.5, the third step follows from the upper bound of $\|B\|_{F}$ and $\|M\|_{F}$.
@@
where the first step follows from Eq. (F.2), the second step follows from Part 2. of Fact F.1, and the third step follows from Part 1. of Fact F.1.
F.3 Sandwich Lower Bound on Frobenius Norm
@@
Lemma F.4 (Formal version of Lemma 5.6).
@@
where the first step, the third step and the fifth step follows from Fact A.4, the second step and the fourth step follows from Fact F.3 and $XX^{\top}\succeq\beta I$.
Taking the square root of both side, we finish the proof. ∎
F.4 Lower Bound on Hadamard Product Between Two Matrices
We present the lemma for lower bound on Hadamard product between two matrices in this section.
Lemma F.5.
@@ F.5 Final Bound
where the second step is from Pythagorean theorem. ∎
@@
Lemma F.7 (Formal version of Lemma 5.7).
@@
F.6 PL Inequality
Here we present the bound for one unit loss function.
Lemma F.8.
@@ F.6 PL Inequality
where the first step follows from Definition B.4 and triangle inequality, the second step follows $x_{1}^{2}+\dots+x_{n}^{2}\leq(x_{1}+\dots+x_{n})^{2}$ when $x_{i}\geq 0$ for any $i\in[n]$. ∎
We present the lemma for proving the PL inequality.
Lemma F.9.
@@
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
  • Let $c(M)$ be defined in Definition B.4.
@@
Lemma F.10 (PL inequality, formal version of 5.5).
@@
  • Let $\mathcal{L}(M)$ be defined in Definition B.7.
@@
We have $\widetilde{f}(M)\cdot{\bf 1}_{n}={\bf 1}_{n}$ and $f\cdot{\bf 1}_{n}={\bf 1}_{n}$ by Definition B.3. Note that $c(M)=\widetilde{f}(M)-f$ by Definition B.4. Thus, we have $c(M)\cdot{\bf 1}_{n}={\bf 0}_{n}$.
@@
where the first and forth steps follow Lemma F.5, the second step follows from Frobenius norm property, the third step follows from triangle inequality, the fifth step follows from Lemma F.8 and Lemma F.9.
@@
where the second step follows from Lemma F.2 and $\alpha_{1}=6\sqrt{n}\max_{i,j\in[d]}\{|W_{i,j}|\}\cdot\|X\|_{F}^{2}\cdot\lambda d$, the third step follows from Lemma F.5 and $\alpha_{2}=\min_{i,j\in[d]}\{|W_{i,j}|\}$, the fourth step follows from Lemma F.4 and $\alpha_{3}=\beta$, the fifth step follows from Lemma F.7 and $\alpha_{4}=\delta$, and the last step follows from $\mu=2\alpha_{2}\cdot\alpha_{3}\cdot\alpha_{4}$ and $\xi=2\alpha_{1}/\mu$.

### v2

@@ D.2 Matrix Form of $\frac{\mathrm{d}}{\mathrm{d}M}{\cal L}(M)$
  • Let ${\cal L}(M)$ be defined in Definition C.7.
  • Let $p$ be defined in Definition D.1.
@@
where the first step follows from Lemma C.15, the second step follows from Lemma D.2, Lemma D.3, and Lemma D.4, the third step follows from basic algebra, and the fourth step follows from Definition D.1. ∎
Appendix E Bounds for Basic Functions
E.1 Basic Assumptions
Here, we introduce our bounded parameters assumption.
Assumption E.1 (Bounded parameters).
@@ E.1 Basic Assumptions
Here, we present the lemma of bounds for $M$ and $M_{c}$.
Lemma E.2 (Bounds for $M$ and $M_{c}$).
@@
This Lemma simply follows from the definition of the Frobenius norm, given that the max value of each entry in $M$ and $M_{c}$ is $1$. ∎
E.2 Bounds for Basic Functions
We first introduce the lemma of bounds for basic functions.
Lemma E.3.
Under Assumption E.1, for all $i_{0}\in[n]$, $j_{0}\in[n]$, $i_{1}\in[d]$, $j_{1}\in[d]$, we have the following bounds
@@ E.2 Bounds for Basic Functions
where the first step follows from Definition C.4, the second step follows the triangle inequality.
@@
Proof of Part 5. The proof simply follows from Assumption E.1 and Fact B.5.
Proof of Part 6. The proof simply follows from Assumption E.1 and Fact B.5.
@@
where the first step follows from Fact B.4 the second step follows from Fact B.5, the third step follows from Part 3., and the last step follows from simple algebra. ∎
E.3 Bounds for Gradient of $\widetilde{f}(M)$
We introduce the lemma of bounds for the gradient of $\widetilde{f}(M)$.
Lemma E.4.
@@ E.3 Bounds for Gradient of $\widetilde{f}(M)$
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
  • Assumption E.1 holds.
@@
Appendix F Lipschitz of Gradient
F.1 Useful Facts
Here, we introduce the fact of the mean value theorem for matrix function.
Fact F.1 (Mean value theorem for matrix function, Fact C.6 in [98]).
@@ F.1 Useful Facts
where the second step follows from Eq. (9), the third step follows from the chain rule, the fourth step follows from the Cauchy-Schwartz inequality.
@@
Here, we introduce the fact of Lipschitz for the product of functions.
Fact F.2 (Lipschitz for product of functions, Fact H.2 in [36]).
@@
  • Let $\{f_{i}(x)\}_{i=1}^{n}$ be a sequence of functions with the same domain and range.
@@ F.2 Lipschitz of $\widetilde{f}(M)$
Lemma F.3 (Lipschitz of $\widetilde{f}(M)$).
@@
  • Assumption E.1 holds.
  • Let $\widetilde{f}(M)$ be defined as Definition C.3.
@@
where the first step follows from Fact F.1, the second step follows from Lemma E.4. ∎
F.3 Lipschitz of $c(M)$
@@
  • Assumption E.1 holds.
  • Let $c(M)$ be defined as Definition C.4.
@@
where the first step follows from Fact F.1, the second step follows from Lemma C.12, the third step follows from Lemma E.4. ∎
F.4 Lipschitz of $\widetilde{f}(M)\circ c(M)$
@@
Lemma F.5 (Lipschitz of $\widetilde{f}(M)\circ c(M)$).
@@
  • Assumption E.1 holds.
  • Let $c(M)$ be defined as Definition C.4.
  • Let $\widetilde{f}(M)$ be defined as Definition C.3.
@@
where the first step follows from triangle inequality, the second step follows from Fact B.5, the third step follows from Lemma E.3, the fourth step follows from Lemma F.4 and Lemma F.3.
@@
F.5 Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})$
@@
Lemma F.6 (Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})$).
@@
  • Assumption E.1 holds.
  • Let $c(M)$ be defined as Definition C.4.
  • Let $\widetilde{f}(M)$ be defined as Definition C.3.
@@
where the first step follows from Fact B.4, the second step follows from basic algebra, the third step follows from Fact B.5, and the fourth step follows from $\|{\bf 1}_{n}\|_{2}=\sqrt{n}$.
@@
which follows from Fact B.5.
Following Eq. (F.5), Eq. (11) and Lemma F.5, we have
@@
F.6 Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})\widetilde{f}(M)$
@@
Lemma F.7 (Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))\cdot{\bf 1}_{n})\widetilde{f}(M)$).
@@
  • Assumption E.1 holds.
  • Let $c(M)$ be defined as Definition C.4.
  • Let $\widetilde{f}(M)$ be defined as Definition C.3.
@@
Following Fact F.2, we have
@@
where we have the upper bound in Lemma E.3, the Lipschitz of $\operatorname{diag}((\widetilde{f}(M)\circ c(M))$ and $\widetilde{f}(M)$ in Lemma F.3 and Lemma F.6. ∎
F.7 Lipschitz of Gradient
@@
Theorem F.8 (Lipschitz of the gradient, formal version of Theorem 5.4).
@@
  • Assumption E.1 holds.
  • Let $c(M)$ be defined as Definition C.4.
  • Let $\widetilde{f}(M)$ be defined as Definition C.3.
@@
where the first step follows from Theorem D.5, and the second step follows the triangle inequality. Now, we proof these two terms separately.
@@
where the first step and the second step follow from Fact B.5, the third step follows from triangle inequality, and the fourth step follows from Assumption E.1.
@@
where the first step follows from Lemma F.5 and Lemma F.7, the second step follows from $n\geq 1$.
@@
which follows from Fact B.5.
@@
which follows from Eq. (F.7), Eq. (F.7), and Eq. (14). ∎
Appendix G Convergence of Gradient Descent
G.1 Helpful Statements
Here, we present useful facts that we use to prove our convergence result.
Fact G.1.
@@ G.1 Helpful Statements
G.2 Lower Bound on Frobenius Norm
In this section, we present the lemma for the lower bound on the Frobenius norm.
Lemma G.2.
@@ G.2 Lower Bound on Frobenius Norm
where the first step follows from Fact B.6, the second step follows from Fact B.5, the third step follows from the upper bound of $\|B\|_{F}$ and $\|M\|_{F}$.
@@
where the first step follows from Eq. (G.2), the second step follows from Part 2. of Fact G.1, and the third step follows from Part 1. of Fact G.1.
G.3 Sandwich Lower Bound on Frobenius Norm
@@
Lemma G.4 (Formal version of Lemma 5.6).
@@
where the first step, the third step, and the fifth step follow from Fact B.4, the second step and the fourth step follows from Fact G.3 and $XX^{\top}\succeq\beta I$.
Taking the square root of both sides, we finish the proof. ∎
G.4 Lower Bound on Hadamard Product Between Two Matrices
We present the lemma for the lower bound on the Hadamard product between two matrices in this section.
Lemma G.5.
@@ G.5 Final Bound
where the second step is from the Pythagorean theorem. ∎
@@
Lemma G.7 (Formal version of Lemma 5.7).
@@
G.6 PL Inequality
Here, we present the bound for one unit loss function.
Lemma G.8.
@@ G.6 PL Inequality
where the first step follows from Definition C.4 and triangle inequality, the second step follows $x_{1}^{2}+\dots+x_{n}^{2}\leq(x_{1}+\dots+x_{n})^{2}$ when $x_{i}\geq 0$ for any $i\in[n]$. ∎
We present the lemma to prove the PL inequality.
Lemma G.9.
@@
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
  • Let $c(M)$ be defined in Definition C.4.
@@
Lemma G.10 (PL inequality, formal version of 5.5).
@@
  • Let $\mathcal{L}(M)$ be defined in Definition C.7.
@@
We have $\widetilde{f}(M)\cdot{\bf 1}_{n}={\bf 1}_{n}$ and $f\cdot{\bf 1}_{n}={\bf 1}_{n}$ by Definition C.3. Note that $c(M)=\widetilde{f}(M)-f$ by Definition C.4. Thus, we have $c(M)\cdot{\bf 1}_{n}={\bf 0}_{n}$.
@@
where the first and fourth steps follow Lemma G.5, the second step follows from Frobenius norm property, the third step follows from triangle inequality, the fifth step follows from Lemma G.8 and Lemma G.9.
@@
where the second step follows from Lemma G.2 and $\alpha_{1}=6\sqrt{n}\max_{i,j\in[d]}\{|W_{i,j}|\}\cdot\|X\|_{F}^{2}\cdot\lambda d$, the third step follows from Lemma G.5 and $\alpha_{2}=\min_{i,j\in[d]}\{|W_{i,j}|\}$, the fourth step follows from Lemma G.4 and $\alpha_{3}=\beta$, the fifth step follows from Lemma G.7 and $\alpha_{4}=\delta$, and the last step follows from $\mu=2\alpha_{2}\cdot\alpha_{3}\cdot\alpha_{4}$ and $\xi=2\alpha_{1}/\mu$. ∎